OpenStudy (anonymous):

How do you figure out X^-4x+y^+8y=-4

Directrix (directrix):

Is the problem written correctly? X^ -4x + y^ +8y = - 4 Are the exponents for x^ and y^ missing?

OpenStudy (anonymous):

No the question is ^ this is means to the second power or square X square - 4x + y square +8y=-4

OpenStudy (tardis):

Oooooh. For future reference, you're supposed to put the exponent next to the ^ If you click right underneath the text reply box, there's an Equation button that would make it a lot clearer. Like so: \[x ^{2}-4x+y ^{2}+8y=-4\]

Directrix (directrix):

>>>How do you figure out What are the specific instructions for the problem? @Cc.j